OpenStudy (anonymous):

Verify that the function satisfies the hypothesis of the Mean Value Theorem on the given interval. Then find all numbers c that satisfy the conclusion of the MVT. f(x) = x/x+2 [1,4]

OpenStudy (amistre64):

the function goes bad at -2; so the interval itself is fine

OpenStudy (amistre64):

as long as there as a slope from end to slope we should be able to satisfy the MVT

OpenStudy (amistre64):

slope from end to end ....

OpenStudy (amistre64):

f(1)=1/3 f(4)=4/6 = 2/3 f(x) moves 1/3, as x moves 3 1/3 --- = 1/6 for a slope 3

OpenStudy (amistre64):

as long as the derivative of the interval has points that match 1/6; we are good

OpenStudy (amistre64):

f(x) = x/x+2 f'(x) = 2 /(x+2) ^2 = 1/6 when (x+2)^2 = 12

OpenStudy (amistre64):

as long as -2+2sqrt(3) is in the interval, wes good

OpenStudy (amistre64):

1.464... looks good

OpenStudy (anonymous):

okay cuz i had a problem wit the derivative

OpenStudy (anonymous):

thank you

OpenStudy (amistre64):

yep